OpenStudy (anonymous):

derivative: d/dx (3 cos (pi t +2)) =?

OpenStudy (anonymous):

I need help just find the derivative using product rule

OpenStudy (anonymous):

i could be wrong

OpenStudy (dumbcow):

not product rule but chain rule \[\frac{d}{dt} 3 \cos(\pi t +2) = -3 \sin(\pi t +2) * \frac{d}{dt}(\pi t +2)\]

OpenStudy (anonymous):

-3sin(pit+2)2

OpenStudy (anonymous):

?

OpenStudy (dumbcow):

no \[d/dt (\pi t +2) = \pi\]

OpenStudy (anonymous):

its not one of the answer choices

OpenStudy (anonymous):

.

OpenStudy (anonymous):

???

OpenStudy (dumbcow):

well first you didn't type it correctly and they want acceleration which is 2nd derivative \[x(t) = 3\cos(\pi t) +2\] \[v(t) = -3 \pi \sin(\pi t)\] \[a(t) = -3 \pi^{2} \cos(\pi t)\]

OpenStudy (anonymous):

thnxs can u help me with another one

OpenStudy (anonymous):

problem 5

OpenStudy (dumbcow):

acceleration is slope of velocity graph.....find where slope is 0 (the max/min points)

OpenStudy (anonymous):

cant tell by the graph?

OpenStudy (dumbcow):

?? just look at turning points